# Knowledge Graphs for the Life Sciences: Recent Developments, Challenges and Opportunities

### v1

@@ 3.1.5. KG storage and inspection
KGs need to be accessible to support a variety of different tasks, beyond the mere integration of different knowledge sources, and thus KG storage management [167, 145, 180] is an active area of research. Current KG storage mechanisms are divided into relation based stores (e.g., [1]) and native graph stores (e.g., [200]). Relational KG stores, either based on relational databases or through NOSQL databases and / or triple stores such as Jena TDB [^15] https://jena.apache.org/documentation/tdb/index.html, have reached a considerable level of maturity and have been optimized in order to avoid common problems, e.g., a large number of null values in columns or optimized query performance [145]. Graph databases store nodes, edges and properties of graphs natively, and support query and graph mining tasks. Examples of state of the art implementations include Neo4J [^16] https://neo4j.com, GraphDB [^17] https://graphdb.ontotext.com, and RDFox [^18] https://www.oxfordsemantic.tech/product. The evolution of the performance of these systems has been the object of systematic studies [9], whereas [171] explicitly focuses on biomedical use cases.
Storage management has implications on the ways KGs support expressive queries for nodes and edges and visualization, to support data analysis, navigation and discovery of related knowledge [96, 165]. Graph databases often provide built-in tools for visualization, e.g., Neo4J, whereas different Javascript libraries (e.g., SigmaJS [^19] https://github.com/jacomyal/sigma.js) are available for developing visualization front ends. Support for complex queries is also either built in a graph database or a triple store by supporting the SPARQL query language [143, 200], or proprietary query languages such as Cypher [49], supported by Neo4J.
3.1.6. Knowledge maintenance and evolution
Given the rapid scientific development in the life sciences, and the consequent continuous update of ontologies for this domain, artefacts annotated with these ontologies can become outdated very quickly, and require some form of update (also called ontology extension). These update mechanisms need to be automated to ensure that they scale to the size of KGs. Automatic update approaches are based on the periodical detection and extraction of new knowledge that is then mapped to existing entities and relations in the KG [186].
@@ 3.1.6. Knowledge maintenance and evolution
Update mechanisms are typically based on the detection of changes [124] that can affect an ontology, e.g., addition, removal or modification of meta-entities (i.e., entities, relations and their definitions). These changes include renaming concepts and properties, setting domain and range restrictions, or setting a subsumption relation. To date, the most comprehensive account of ontology change is given in [48], where change is described for different sub-fields, e.g., ontology alignment, matching and mapping, morphisms, articulation, translation, evolution, debugging, versioning, integration and merging; each with different requirements and implications. The study [140] further investigates the impact of biomedical ontology evolution on materialization.
Currently available tools and methodologies use (semi)-automated methods to perform many of the operations that trigger a change in an ontology and the consequent creation of a new version [56, 65]. Different ontology management platforms and portals mandate different principles and frameworks for handling ontology versioning (e.g., OBO foundry [^20] http://www.obofoundry.org/principles/fp-004-versioning.html or BioPortal [^21] https://bioportal.bioontology.org), but these are typically implemented by ontology developers with limited tool support. Section 3.3 presents an example of automated ontology extension that relies on machine learning to cope with the scale of data.
3.2. Examples of Life Science KG Construction
In this section we provide two examples of life science KGs that illustrate in practice the phases composing the generic KG construction process discussed in Section 3; namely a KG for Pharmacogenomics, PGxLOD [121], and one for Ecotoxicological Analysis, TERA [127, 128].
@@ 3.2. Examples of Life Science KG Construction
Alignment for Knowledge Validation: An Example of Pharmacogenomics. As mentioned in Section 3, the task of aligning knowledge in KGs supports several downstream applications and domains. For instance, pharmacogenomics studies the influence of genetic factors on drug response phenotypes (e.g., expected effect, side effect). Hence, pharmacogenomics is of interest for personalized medicine. The atomic knowledge unit in pharmacogenomics is a ternary relationship between a drug, a genetic factor, and a phenotype. Such a relationship states that a patient being treated with the specified drug while having the specified genetic factor may experience the described phenotype. Semantic Web and KG technologies have been employed in this application domain, for example by building ontologies in which patients and pharmacogenomic knowledge are represented, and then using deductive reasoning mechanism to conditionally recommend genetic testing before drug prescription [156]. However, the knowledge relevant to pharmacogenomics is scattered across several sources including reference databases such as PharmGKB, and the biomedical literature. Additionally, this knowledge may lack sufficient validation to be implemented in clinical practice. For example, some relationships may have only been observed in smaller cohorts of patients or in non-replicated studies. Hence, there is a need to align different sources of pharmacogenomic knowledge to detect additional evidence validating (or moderating) a knowledge unit. To this aim, the PGxLOD KG was proposed [121]. Automatic knowledge extraction approaches were applied on semi-structured and unstructured data from PharmGKB and the biomedical literature to represent their knowledge in the KG. Then, matching approaches were developed to align knowledge units from various sources [120, 122]. The resulting alignments outlined some agreements between PharmGKB and the biomedical literature, which was expected since PharmGKB is manually completed by experts after reviewing the literature. Interestingly, this automatic knowledge extraction pipeline could guide the manual review process by automatically pointing out studies confirming or mentioning a pharmacogenomic knowledge unit.
Knowledge Integration: An Example of Ecotoxicological Analysis. In ecotoxicological analysis, data and knowledge from different domains such as chemistry and biology are often needed. These are usually located in different sources such as spreadsheets or CSV files for local experimental results, open databases for public research results, and ontologies for domain knowledge. Thus knowledge integration becomes a critical and fundamental challenge before real analysis can be conducted. In the study by Myklebust et al. [127, 128], which aims to predict adverse biological effects of chemicals on species, a toxicological effect and risk assessment KG named TERA was constructed for knowledge integration. TERA includes three sub-KGs: (i) the Chemical sub-KG, which is constructed by integrating the vocabulary MeSH (Medical Subject Headings) with selective knowledge from two chemical databases PubChem and ChEMBL utilizing the chemical mappings in Wikidata; (ii) the Taxonomy sub-KG, which is constructed by integrating EOL (Environment Ontology for Livestock) and the NCBITaxon ontology utilizing NIBI-EOL mappings in Wikidata; and (iii) the ECOTOX sub-KG, which is composed of RDF triples transformed from experimental risk results and is aligned with the other two sub-KGs by the ontology alignment system LogMap [82] and the chemical mappings in Wikidata. Another example of knowledge integration is for drug repurposing, where the KG Hetionet [^22] https://github.com/hetio/hetionet is created by integrating 29 public resources, including biomedical KGs and other types of data [69].
3.3. What has been done: recent technical developments
Given the many existing ontologies in life sciences, e.g., ontologies available in the OBO Foundry collection or in BioPortal [135], KG construction usually involves the reuse, alignment, and enrichment of state-of-the-art ontologies. The existing ontologies in life sciences need to be updated given the new discoveries in the field. This is broadly a key issue in the management, maintenance, and evolution of ontologies. We select a few promising use cases below to highlight some recent developments that support the KG construction in the life sciences.
@@ 3.3. What has been done: recent technical developments
Repositories of Ontologies and Mappings. Ontologies and their mappings play a central role in semantically enabled products and services consumed by life science companies, academic institutions and universities, as highlighted by the Pistoia Alliance ontology mapping project [60]. [^23] https://www.pistoiaalliance.org/projects/current-projects/ontologies-mapping/ Ontology mappings are essential in knowledge graph construction tasks to bridge the knowledge provided by different ontologies and expand their coverage. Ontology mappings can also play a key role when identifying the right ontologies to be reused as they will enable the retrieval of the relevant (overlapping) ontologies for the domain of interest. For this reason, a number of notable efforts in life sciences have created large repositories of ontologies and mappings to serve the research within the community. Prominent examples include the UMLS Metathesaurus [10], BioPortal [135, 155], MONDO [175], and the EBI services: OLS [177], OxO [86] and the RDF platform [87]. The UMLS Metathesaurus is a comprehensive effort for integrating biomedical ontologies through mappings. In its 2023AA version, it integrates more than two hundred vocabularies, with more than 3 million unique concepts and more than 15 million concept names. BioPortal is a repository containing more than 1,000 biomedical ontologies and more than 79 million lexically computed mappings among them (as of July 13, 2023). The Mondo Disease Ontology (MONDO) is a manually curated effort to harmonize and integrate disease conceptualizations and definitions across state-of-the-art ontologies (e.g., HPO [99], DO [158], ICD, SNOMED CT, etc.). The services provided by the European Bioinformatics Institute (EBI) also deserve a special mention. The Ontology Lookup Service (OLS) has become a reference to explore the latest versions of more than two hundred ontologies via its graphical interface or programmatically via its API. OxO is a repository of ontology mappings and cross-references extracted from the OLS and UMLS. OxO allows users to visually traverse the graph of mappings to identify additional potential mappings beyond direct ones (i.e., multi-hop mappings). Finally, the EBI RDF platform provides a unified KG with all the RDF resources at the EBI. Complementary to the efforts from the life sciences, the Semantic Web has also contributed to the systematic evaluation of mappings in public repositories (e.g., [83, 46]) and mappings produced by automated ontology mapping systems (e.g., the Ontology Alignment Evaluation Initiative (OAEI) [141]). Automatically generated mappings of high quality have the potential to be integrated within the aforementioned repositories and hence, the OAEI has always had a special focus on life science test cases with evaluation tracks like Anatomy [41], LargeBio [85], Phenotype [61] and the newly created track BioML [66]. The Simple Standard for Sharing Ontological Mappings (SSSOM) [117] represents a joint effort between the life sciences and Semantic Web communities to facilitate the exchange of mappings across different parties and repositories, while keeping the provenance and other relevant characteristics of the mappings.
Ontology Extension. Ontology extension in life sciences aims to connect new concepts and their relations to an ontology from updated sources, e.g., scientific papers in PubMed and chemical information in PubChem [^24] https://pubchem.ncbi.nlm.nih.gov/. Manual ontology extension, while essential for the development of gold standard resources, is not scalable to the full scope of large domains due to its high cost and low efficiency, and sometimes is even unfeasible as human beings may not be able to review the quantities of new information at the rate they become available. Thus machine-learning-based, automated methods are needed. One recent example is the use of deep learning, specifically a Transformer-based model, to categorize new chemical entities within the ChEBI ontology [^25] https://www.ebi.ac.uk/chebi/ [55]. In addition, recent studies have explored enriching SNOMED CT by mining new concepts from texts [36] and placing them into the ontology [112, 35]. A new concept can be identified by NIL entity linking, i.e., exploring unlinkable mentions, usually through setting a “linkable” score threshold or through classification [36]. Resolution and disambiguation of NIL mentions with clustering can help to represent NIL entities [68, 94]. For concept placement, similar to the aforementioned CHEBI ontology extension [55], machine learning, especially in the form of Transformer-based deep learning, has been applied to predict subsumption relations between a new concept and the existing concepts. Complex concepts in OWL ontologies that contain logical operators (e.g., existential quantifier and conjunction in SNOMED CT) can be supported in subsumption prediction [24] and new concept placement [35]. Another group of studies use post-coordination or formalising a new term with existing concepts and attributes [17, 95], which is similar to composing subsumption axioms with complex concepts. The methods include using lexical features [95], word embeddings and KG embeddings [17]. Pre-trained and Large Language Models, through fine-tuning, zero-shot and few-shot prompting have the potential to support the mining [36] and placement of new concepts (e.g., by subsumption prediction [24, 67]).
Instance Matching: Automated Clinical Coding. A main source for patients’ KG construction is Electrical Health Records (EHR). Using medical ontologies as backbones, it is possible to add a layer of data by instance matching (or patient matching) through Clinical Coding. Clinical coding is the task of transforming medical information in EHR into structured codes described in medical ontologies [37], e.g., ICD and SNOMED CT. Recent approaches mainly formulate the problem as a multi-label classification problem. Various neural network architectures have been proposed and knowledge plays a key role to enhance the neural architectures [37, 81]. Pre-trained language models, e.g., BERT [33], have been applied to clinical coding and gradually achieved better results with adapted modelling methods and more advanced language models, e.g., PLM-ICD [72] with RoBERTa-PM [107], according to studies [37, 44, 80]. Other studies formulate the task as a Named Entity Recognition and Linking (NER+L) problem, by extraction of concepts and linking them with the ontologies [37]. Overall, the recent progress in clinical coding, along with the advent of Large Language Models (LLMs) suggests a trend in this area for patients’ KG construction from EHR. However, there is still room for improvement in knowledge integration to better address explainability (see Section 5 for more details) and in zero-shot learning problems, i.e., for classifying into rare codes or concepts [37, 44, 81]. There are also further recent examples of instance matching with EHR data, including the works [16, 169].
3.4. What are the challenges?
@@
4. Life Science Knowledge Discovery
Research into AI technologies – including machine learning and KG-based reasoning – to accelerate the pace of scientific discovery is an emerging and rapidly developing field. The challenge lies in assisting scientists to uncover new knowledge and solutions, such as discovering novel therapeutic opportunities, identifying candidate molecular drugs to treat complex diseases or alternatively new uses for existing drugs, and supporting more personalized predictions.

### v4

@@ 3.1.5 KG storage and inspection
KGs need to be accessible to support a variety of different tasks, beyond the mere integration of different knowledge sources, and thus KG storage management [164, 142, 177] is an active area of research. Current KG storage mechanisms are divided into relation based stores (e.g., [1]) and native graph stores (e.g., [197]). Relational KG stores, either based on relational databases or through NOSQL databases and / or triple stores such as Jena TDB [^12] https://jena.apache.org/documentation/tdb/index.html, have reached a considerable level of maturity and have been optimized in order to avoid common problems, e.g., a large number of null values in columns or optimized query performance [142]. Graph databases store nodes, edges and properties of graphs natively, and support query and graph mining tasks. Examples of state of the art implementations include Neo4J [^13] https://neo4j.com, GraphDB [^14] https://graphdb.ontotext.com, and RDFox [^15] https://www.oxfordsemantic.tech/product. The evolution of the performance of these systems has been the object of systematic studies [8], whereas [168] explicitly focuses on biomedical use cases. Storage management has implications on the ways KGs support expressive queries for nodes and edges and visualization, to support data analysis, navigation and discovery of related knowledge [93, 162]. Graph databases often provide built-in tools for visualization, e.g., Neo4J, whereas different Javascript libraries (e.g., SigmaJS [^16] https://github.com/jacomyal/sigma.js) are available for developing visualization front ends. Support for complex queries is also either built in a graph database or a triple store by supporting the SPARQL query language [140, 197], or proprietary query languages such as Cypher [47], supported by Neo4J.
3.1.6 Knowledge maintenance and evolution
@@
Given the rapid scientific development in the life sciences, and the consequent continuous update of ontologies for this domain, artifacts annotated with these ontologies can become outdated very quickly, and require some form of update (also called ontology extension). These update mechanisms need to be automated to ensure that they scale to the size of KGs. Automatic update approaches are based on the periodical detection and extraction of new knowledge that is then mapped to existing entities and relations in the KG [183]. Update mechanisms are typically based on the detection of changes [121] that can affect an ontology, e.g., addition, removal or modification of meta-entities (i.e., entities, relations and their definitions). These changes include renaming concepts and properties, setting domain and range restrictions, or setting a subsumption relation. To date, the most comprehensive account of ontology change is given in [46], where change is described for different sub-fields, e.g., ontology alignment, matching and mapping, morphisms, articulation, translation, evolution, debugging, versioning, integration and merging; each with different requirements and implications. The study [137] further investigates the impact of biomedical ontology evolution on materialization. Currently available tools and methodologies use (semi)-automated methods to perform many of the operations that trigger a change in an ontology and the consequent creation of a new version [54, 63]. Different ontology management platforms and portals mandate different principles and frameworks for handling ontology versioning (e.g., OBO foundry [^17] http://www.obofoundry.org/principles/fp-004-versioning.html or BioPortal [^18] https://bioportal.bioontology.org), but these are typically implemented by ontology developers with limited tool support. Section 3.3 presents an example of automated ontology extension that relies on machine learning to cope with the scale of data.
3.2 Examples of Life Science KG Construction
@@
In this section we provide two examples of life science KGs that illustrate in practice the phases composing the generic KG construction process discussed in Section 3; namely a KG for Pharmacogenomics, PGxLOD [118], and one for Ecotoxicological Analysis, TERA [124, 125]. Alignment for Knowledge Validation: An Example of Pharmacogenomics. As mentioned in Section 3, the task of aligning knowledge in KGs supports several downstream applications and domains. For instance, pharmacogenomics studies the influence of genetic factors on drug response phenotypes (e.g., expected effect, side effect). Hence, pharmacogenomics is of interest for personalized medicine. The atomic knowledge unit in pharmacogenomics is a ternary relationship between a drug, a genetic factor, and a phenotype. Such a relationship states that a patient being treated with the specified drug while having the specified genetic factor may experience the described phenotype. Semantic Web and KG technologies have been employed in this application domain, for example by building ontologies in which patients and pharmacogenomic knowledge are represented, and then using deductive reasoning mechanisms to conditionally recommend genetic testing before drug prescription [153]. However, the knowledge relevant to pharmacogenomics is scattered across several sources including reference databases such as PharmGKB, and the biomedical literature. Additionally, this knowledge may lack sufficient validation to be implemented in clinical practice. For example, some relationships may have only been observed in smaller cohorts of patients or in non-replicated studies. Hence, there is a need to align different sources of pharmacogenomic knowledge to detect additional evidence validating (or moderating) a knowledge unit. To this aim, the PGxLOD KG was proposed [118]. Automatic knowledge extraction approaches were applied on semi-structured and unstructured data from PharmGKB and the biomedical literature to represent their knowledge in the KG. Then, matching approaches were developed to align knowledge units from various sources [117, 119]. The resulting alignments outlined some agreements between PharmGKB and the biomedical literature, which was expected since PharmGKB is manually completed by experts after reviewing the literature. Interestingly, this automatic knowledge extraction pipeline could guide the manual review process by automatically pointing out studies confirming or mentioning a pharmacogenomic knowledge unit. Knowledge Integration: An Example of Ecotoxicological Analysis. In ecotoxicological analysis, data and knowledge from different domains such as chemistry and biology are often needed. These are usually located in different sources such as spreadsheets or CSV files for local experimental results, open databases for public research results, and ontologies for domain knowledge. Thus knowledge integration becomes a critical and fundamental challenge before real analysis can be conducted. In the study by Myklebust et al. [124, 125], which aims to predict adverse biological effects of chemicals on species, a toxicological effect and risk assessment KG named TERA was constructed for knowledge integration. TERA includes three sub-KGs: (i) the Chemical sub-KG, which is constructed by integrating the vocabulary MeSH (Medical Subject Headings) with selective knowledge from two chemical databases PubChem and ChEMBL utilizing the chemical mappings in Wikidata; (ii) the Taxonomy sub-KG, which is constructed by integrating EOL (Environment Ontology for Livestock) and the NCBITaxon ontology utilizing NIBI-EOL mappings in Wikidata; and (iii) the ECOTOX sub-KG, which is composed of RDF triples transformed from experimental risk results and is aligned with the other two sub-KGs by the ontology alignment system LogMap [80] and the chemical mappings in Wikidata. Another example of knowledge integration is for drug repurposing, where the KG Hetionet [^19] https://github.com/hetio/hetionet is created by integrating 29 public resources, including biomedical KGs and other types of data [67].
3.3 What has been done: recent technical developments
@@
Given the many existing ontologies in life sciences, e.g., ontologies available in the OBO Foundry collection or in BioPortal [132], KG construction usually involves the reuse, alignment, and enrichment of state-of-the-art ontologies. The existing ontologies in life sciences need to be updated given the new discoveries in the field. This is broadly a key issue in the management, maintenance, and evolution of ontologies. We select a few promising use cases below to highlight some recent developments that support the KG construction in the life sciences. Repositories of Ontologies and Mappings. Ontologies and their mappings play a central role in semantically enabled products and services consumed by life science companies, academic institutions and universities, as highlighted by the Pistoia Alliance ontology mapping project [58]. [^20] https://www.pistoiaalliance.org/projects/current-projects/ontologies-mapping/ Ontology mappings are essential in knowledge graph construction tasks to bridge the knowledge provided by different ontologies and expand their coverage. Ontology mappings can also play a key role when identifying the right ontologies to be reused as they will enable the retrieval of the relevant (overlapping) ontologies for the domain of interest. For this reason, a number of notable efforts in life sciences have created large repositories of ontologies and mappings to serve the research within the community. Prominent examples include the UMLS Metathesaurus [9], BioPortal [132, 152], MONDO [172], and the EBI services: OLS [174], OxO [84] and the RDF platform [85]. The UMLS Metathesaurus is a comprehensive effort for integrating biomedical ontologies through mappings. In its 2023AA version, it integrates more than two hundred vocabularies, with more than 3 million unique concepts and more than 15 million concept names. BioPortal is a repository containing more than 1,000 biomedical ontologies and more than 79 million lexically computed mappings among them (as of July 13, 2023). The Mondo Disease Ontology (MONDO) is a manually curated effort to harmonize and integrate disease conceptualizations and definitions across state-of-the-art ontologies (e.g., HPO [96], DO [155], ICD, SNOMED CT, etc.). The services provided by the European Bioinformatics Institute (EBI) also deserve a special mention. The Ontology Lookup Service (OLS) has become a reference to explore the latest versions of more than two hundred ontologies via its graphical interface or programmatically via its API. OxO is a repository of ontology mappings and cross-references extracted from the OLS and UMLS. OxO allows users to visually traverse the graph of mappings to identify additional potential mappings beyond direct ones (i.e., multi-hop mappings). Finally, the EBI RDF platform provides a unified KG with all the RDF resources at the EBI. Complementary to the efforts from the life sciences, the Semantic Web has also contributed to the systematic evaluation of mappings in public repositories (e.g., [81, 44]) and mappings produced by automated ontology mapping systems (e.g., the Ontology Alignment Evaluation Initiative (OAEI) [138]). Automatically generated mappings of high quality have the potential to be integrated within the aforementioned repositories and hence, the OAEI has always had a special focus on life science test cases with evaluation tracks like Anatomy [39], LargeBio [83], Phenotype [59] and the newly created track BioML [64]. The Simple Standard for Sharing Ontological Mappings (SSSOM) [114] represents a joint effort between the life sciences and Semantic Web communities to facilitate the exchange of mappings across different parties and repositories, while keeping the provenance and other relevant characteristics of the mappings. Ontology Extension. Ontology extension in life sciences aims to connect new concepts and their relations to an ontology from updated sources, e.g., scientific papers in PubMed and chemical information in PubChem [^21] https://pubchem.ncbi.nlm.nih.gov/. Manual ontology extension, while essential for the development of gold standard resources, is not scalable to the full scope of large domains due to its high cost and low efficiency, and sometimes is even unfeasible as human beings may not be able to review the quantities of new information at the rate they become available. Thus machine-learning-based, automated methods are needed. One recent example is the use of deep learning, specifically a Transformer-based model, to categorize new chemical entities within the ChEBI ontology [^22] https://www.ebi.ac.uk/chebi/ [53]. In addition, recent studies have explored enriching SNOMED CT by mining new concepts from texts [35] and placing them into the ontology [109, 34]. A new concept can be identified by NIL entity linking, i.e., exploring unlinkable mentions, usually through setting a “linkable” score threshold or through classification [35]. Resolution and disambiguation of NIL mentions with clustering can help to represent NIL entities [66, 91]. For concept placement, similar to the aforementioned CHEBI ontology extension [53], machine learning, especially in the form of Transformer-based deep learning, has been applied to predict subsumption relations between a new concept and the existing concepts. Complex concepts in OWL ontologies that contain logical operators (e.g., existential quantifier and conjunction in SNOMED CT) can be supported in subsumption prediction [23] and new concept placement [34]. Another group of studies use post-coordination or formalising a new term with existing concepts and attributes [16, 92], which is similar to composing subsumption axioms with complex concepts. The methods include using lexical features [92], word embeddings and KG embeddings [16]. Pre-trained and Large Language Models, through fine-tuning, zero-shot and few-shot prompting have the potential to support the mining [35] and placement of new concepts (e.g., by subsumption prediction [23, 65]). Instance Matching: Automated Clinical Coding. A main source for patients’ KG construction is Electronic Health Records (EHR). Using medical ontologies as backbones, it is possible to add a layer of data by instance matching (or patient matching) through Clinical Coding. Clinical coding is the task of transforming medical information in EHR into structured codes described in medical ontologies [36], e.g., ICD and SNOMED CT. Recent approaches mainly formulate the problem as a multi-label classification problem. Various neural network architectures have been proposed and knowledge plays a key role to enhance the neural architectures [36, 79]. Pre-trained language models, e.g., BERT [32], have been applied to clinical coding and gradually achieved better results with adapted modeling methods and more advanced language models, e.g., PLM-ICD [70] with RoBERTa-PM [104], according to studies [36, 42, 78]. Other studies formulate the task as a Named Entity Recognition and Linking (NER+L) problem, by extraction of concepts and linking them with the ontologies [36]. Overall, the recent progress in clinical coding, along with the advent of Large Language Models (LLMs) suggests a trend in this area for patients’ KG construction from EHR. However, there is still room for improvement in knowledge integration to better address explainability (see Section 5 for more details) and in zero-shot learning problems, i.e., for classifying into rare codes or concepts [36, 42, 79]. There are also further recent examples of instance matching with EHR data, including the works [15, 166].
3.4 What are the challenges?
@@
4 Life Science Knowledge Discovery